You are riding your bicycle directly away from a stationary source of sound and hear a frequency that is 1.4% lower than the emitted frequency. If the speed of sound is 344 m/s, what is your speed?
1
Expert's answer
2018-03-16T08:05:08-0400
f = [(V-V_r)/(V+V_s )] f_o
A stationary source of sound:
V_s=0
f/f_0 =1-V_r/V
V_r/V=1-f/f_0 =0.014
V_r=0.014(344)=4.82 m/s.